gary has a total of 27 dvd's and cd's. the number of dvd's is 3 less than twice the number of cd's. how many of each does he have?
Answer by oberobic(2304)

x = number of dvd
y = number of cd
x + y = 27
x = 2y - 3
.
set up a simultaneous equations
x + y = 27
x - 2y = -3
subtract
3y = 30
y = 10
.
so
x = 27 - 10 = 17
.
Answer:
10 CDs
17 DVDs
.
Done